White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates: Department Cutting Finishing Direct labor-hours 8,700 86,000 Machine-hours 69,100 2,400 Total fixed manufacturing overhead cost $ 360,000 $ 407,000 Variable manufacturing overhead per machine-hour $ 3.00 — Variable manufacturing overhead per direct labor-hour — $ 3.75 Required: 1. Compute the predetermined overhead rate for each department. 2. The job cost sheet for Job 203, which was started and completed during the year, showed the following: Department Cutting Finishing Direct labor-hours 5 12 Machine-hours 89 5 Direct materials $ 710 $ 370 Direct labor cost $ 115 $ 276 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203. 3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates?
Answer #1

1)Predetermined overhead rate =Estimated overhead /estimated activity

where activity can be machine hours or direct labor hours

Cutting Finishing
variable overhead rate 3 3.75
Fixed overhead rate 360000/69100= $ 5.21 (rounded) 407000/86000 =$ 4.73
Total Predetermined overhead rate $ 8.21 per machine hours $ 8.48 per direct labor hours

2)

Cutting Finishing Total
Direct material 710 370 1080
Direct labor 115 276 391
overhead applied 89*8.21=730.69 12*8.48=101.76 832.45
Total cost 1555.69 747.76 2303.45

3/)

If company uses plantwide Predetermined overhead rate =Estimated total overhead /Estimated total direct labor hours

                = 1296800/94700

                = $ 13.69 per direct labor hours

Overhead assigned to Job 203 using plantwide overhead rate =total actual direct labor hours *overhead rate

           = (5+12)*13.69

              = $ 232.73

yes ,if a company uses plantwide overhead rate(232.73) then ,overhead amount will be substantially different from departmental overhead rate (832.45)

working :

Estimated total overhead is calculated as follows :

Cutting Finishing
variable overhead 3*69100=207300 3.75*86000=322500
Fixed overhead 360000 407000
Total estimated overhead 567300 729500

Total estimated overhead = 567300+729500= 1296800

Total estimated direct labor hours= 8700+86000=94700